OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! A ball is thrown vertically upward from the top of a 96‐ft tower, with an initial velocity of 16 ft/sec. Its position function is s(t) = –16t2 + 16t + 96. What is its velocity in ft/sec when it hits the ground? a. -32 b. -48 c. -60 d. -80

jimthompson5910 (jim_thompson5910):

is this for a calculus class?

OpenStudy (anonymous):

How do you get from a position function to a velocity function?

OpenStudy (aryana_maria2323):

Yes calculus honors.

OpenStudy (aryana_maria2323):

That I do not know

jimthompson5910 (jim_thompson5910):

I'm assuming you've learned derivatives yet? or that has yet to come up in the lesson?

OpenStudy (anonymous):

To get from a position function to a velocity function, you take the derivative of the velocity function. if s(t) is position and v(t) is velocity, then s'(t)=v(t) Can you take the derivative of s(t) for me?

OpenStudy (aryana_maria2323):

Yes I just learned about derivatives.

jimthompson5910 (jim_thompson5910):

so if s(t) = –16t^2 + 16t + 96, then what is v(t) ?

OpenStudy (aryana_maria2323):

The derivative of s(t) is -32t+16

jimthompson5910 (jim_thompson5910):

so v(t) = -32t+16

jimthompson5910 (jim_thompson5910):

if you knew when the object the ground, you can figure out its speed to figure out when the object hits the ground, solve s(t) = 0 for t

jimthompson5910 (jim_thompson5910):

s(t) = –16t^2 + 16t + 96 0 = –16t^2 + 16t + 96 ... plug in s(t) = 0 use the quadratic formula to solve for t t =??? or t = ???

OpenStudy (aryana_maria2323):

I got t=-2,3 or I did it another way and replaced all the "t's" with 0 and got 96.

OpenStudy (aryana_maria2323):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`replaced all the "t's" with 0 and got 96` that's if you wanted the y intercept. We want the x intercepts in this case

jimthompson5910 (jim_thompson5910):

t = -2 or t = 3 are the correct solutions to s(t) = 0

jimthompson5910 (jim_thompson5910):

ignore the negative solution because t = -2 makes no sense. You cannot have negative time values

OpenStudy (aryana_maria2323):

okay so the first way I did correct then. t= -2 or t=3. So then what do I do?

OpenStudy (aryana_maria2323):

Then why are all the answer choices - numbers?

jimthompson5910 (jim_thompson5910):

the only practical t value is t = 3 the rocket will hit the ground at t = 3 now plug t = 3 into the v(t) equation

jimthompson5910 (jim_thompson5910):

`Then why are all the answer choices - numbers? ` they are negative to indicate negative velocities positive velocities are when the object is moving up negative velocities are when the object is moving down

OpenStudy (aryana_maria2323):

I got t=1/2 - 3v/32

jimthompson5910 (jim_thompson5910):

v(t) = -32t + 16 v(3) = -32(3) + 16 ... plug in t = 3 v(3) = ______________ (fill in the blank)

OpenStudy (aryana_maria2323):

-80/3

OpenStudy (aryana_maria2323):

sorry -80

OpenStudy (aryana_maria2323):

If that is correct can you help me with one more problem? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-80 is the correct answer

jimthompson5910 (jim_thompson5910):

ok make a new post and tag me in it

OpenStudy (aryana_maria2323):

Okay thank you.